# Cameras as Relative Positional Encoding

### v1

@@ 3.4 Projective Position Encoding (PRoPE)
To implement PRoPE, we define a new set of $\mathbf{D}^{\texttt{PRoPE}}_{t}\in\mathbb{R}^{d\times d}$ matrices and use GTA-style attention (Equation 15) to inject them into transformer blocks. We design these matrices to (1) encode frustum relationships between cameras—this uses the projective relationship in Equation 17—and (2) encode relative patch positions within cameras—this follows GTA [3] and uses RoPE terms. These goals are achieved with complementary submatrices, each with shape $\frac{d}{2}\times\frac{d}{2}$:
@@
In these definitions, $\text{RoPE}_{d/4}(\cdot)$ constructs $\frac{d}{4}\times\frac{d}{4}$ rotary embeddings [1] for $(x_{t},y_{t})$, which are the patch coordinates for token $t$.
Properties.
@@
  3. 3.
    Reduction to RoPE [28]. Equation 21 evaluates to identity for patches from the same image. For these token pairs, PRoPE reduces to the remaining terms in $\mathbf{D}_{t}^{\texttt{PRoPE}}$: the RoPE terms used by single-image vision transformers.
@@ 4 Experiments
We structure our investigation as follows. We begin with basic experimental setup (Section 4.1), followed by empirical comparison of absolute and relative encodings of camera geometry (Section 4.2). We then evaluate PRoPE’s ability to understand camera intrinsics in more detail (Section 4.3), followed by an exploration of hybrid approaches that combine both attention-level and token-level conditioning (Section 4.4). Finally, we present studies on generalization: to out-of-distribution test inputs (Section 4.5), with additional tasks (Section 4.6), and when scaling to larger models and more compute (Section 4.7).
@@ 4.1 Experiment Setup
We include metrics for several camera conditioning techniques—Näive and Plücker raymap encodings, CAPE [4], GTA [3], and PRoPE. In our experiments, GTA refers to the SE(3)+SO(2) variation studied by [3], where SO(2) refers to RoPE on patch positions. As discussed in Section 3.4, PRoPE adopts the self-attention mechanism and RoPE combination proposed by GTA. The primary difference between PRoPE and GTA is therefore the use of relative projective relationships instead of relative SE(3) relationships between cameras.
While we also include results on additional tasks (Section 4.6), iur core experiments evaluate camera conditioning techniques using feedforward novel view synthesis (NVS). NVS is an ideal benchmarking task because it requires fine-grained geometric reasoning: models are trained to render scenes from target viewpoints, given only calibrated reference images and target camera parameters. We do this by reimplementing and training variants of LVSM [8], a state-of-the-art novel view synthesis method that originally encoded camera geometry with Plücker raymaps. We train and evaluate separately on the RealEstate10K [12] and Objaverse [13] datasets. Each scene in RealEstate10K is captured with the same camera (i.e., constant intrinsics), but cameras can vary between scenes. The Objaverse renders we use, on the other hand, use the same camera intrinsics across the entire dataset.
We take several steps to ensure fairness in evaluations. Models are trained in the same codebase, with matched hyperparameters and optimization steps. We pad images with a fixed embedding when raymaps are not used as input (CAPE [4], GTA [3], and PRoPE); this lets all experiments use identical input, output, and overall model sizes. Finally, while our initial experiments use smaller model sizes ($\sim$25M parameters), we also validate larger ones in Section 4.7. More details are provided in Appendix A.1.1.
@@ 4.2 Relative vs Absolute Positional Encodings
Method RealEstate10K [12] Objaverse [13]
@@
Näive Raymap 20.54 0.210 0.623 21.59 0.153 0.856
CAPE [4] 21.11 0.234 0.656 19.68 0.220 0.827
GTA [3] 22.51 0.164 0.707 23.70 0.104 0.879
@@
Table 1: Novel view synthesis comparison, with constant intrinsics in each scene. We compare different camera conditioning approaches applied to the LVSM [8] framework.
Results for each of the four encoding techniques on both datasets are presented in Table 1 and discussed below.
Relative encodings outperform absolute ones. Consistent with prior work [4, 3], we observe that relative encodings for camera geometry consistently outperform absolute ones. CAPE, GTA, and PRoPE all yield improvements over the widely used Plücker raymaps, with PRoPE (followed closely by GTA) producing the best results.
Projective positional encoding improves view synthesis quality. PRoPE consistently outperforms other encoding methods across metrics on RealEstate10K [12], despite the dataset’s limited intrinsics variation. This confirms that capturing more complete camera information (both intrinsic and extrinsic) in our relative encoding is beneficial. We also find no loss of performance when the train and test images all have constant intrinsics: GTA [3] and PRoPE produce identical metrics for the Objaverse dataset, which further verifies that PRoPE reduces to GTA when model does not need to care about camera intrinsics.
Method RealEstate10K [12] Objaverse [13]
@@
Naïve Raymap 20.56 0.301 0.629 21.00 0.191 0.845
CAPE [4] 15.94 0.497 0.699 16.78 0.322 0.760
GTA [3] 15.77 0.512 0.641 18.00 0.257 0.775
@@
Table 2: Novel view synthesis, with varying intrinsics in each scene. We compare the LVSM [8] model trained with different camera condition strategies, on intrinsics-augmented dataset variants.
@@ 4.3 Attention-Level Intrinsics Conditioning
Real-world data often involves different cameras and focal lengths—consider multiview rigs on autonomous cars or zoom lenses on point-and-shoot cameras. The ability to understand intrinsics is the primary difference between PRoPE’s relative camera encoding and prior relative pose encoding techniques. Existing datasets such as RealEstate10K [12], however, typically simulate multiview scenarios using a fixed camera setup with identical intrinsics across all viewpoints.
To understand how effective PRoPE is at encoding intrinsics information, we evaluate each conditioning method on intrinsics-augmented versions of the RealEstate10K [12] and Objaverse [13] datasets. We augment RealEstate10K by applying a zoom factor sampled uniformly in $[1,3]$ to each image. For Objaverse, we switch from constant field of views to uniformly sampled ones between 35 and 50 degrees. In contrast to Section 4.2, this means that cameras within scenes can vary in both extrinsics and intrinsics. We present quantitative results in Table 2, qualitative results in Figure A.1, and findings below.
PRoPE enables intrinsic-aware multiview understanding. We observe that PRoPE outperforms all alternative camera conditioning techniques for both datasets. Existing attention-based methods, which let networks condition on relative pose, perform poorly without knowledge of intrinsics. While token-level raymaps carry sufficient camera information, they perform worse overall than PRoPE’s relative conditioning formulation.
@@ 4.4 Hybrid Encoding Strategies
Figure 2: Hybrid camera encoding. Left: token-level conditioning using only raymaps, which capture both camera intrinsics and extrinsics. Right: hybrid encoding where the attention-level PRoPE encoding captures relationships between camera frustums, and token-level CamRay raymaps encode local ray directions.
Method RealEstate10K [12] Objaverse [13]
@@
Table 3: Novel view synthesis with hybrid camera encodings, with varying intrinsics in each scene. Intrinsics conditioning can be added to the SE(3) GTA encoding by concatenating local frame camera rays to the network input.
Token- and attention-level camera encodings require modifications to differents parts of the transformer architecture. They are therefore compatible with each other: both conditioning styles can be used simultaneously. To compare PRoPE against a stronger baseline while simultaneously exploring these “hybrid” conditioning strategies (Figure 2), we train LVSM variations that couple relative encodings with local, camera-frame raymaps:
$\mathbf{M}_{i,CamRay{}}^{u,v}$ $=\mathbf{R}^{\text{cw}}_{i}\mathbf{d}_{i}^{u,v}\propto\boldsymbol{K}_{i}^{-1}\begin{bmatrix}u&v&1\end{bmatrix}^{\top}\in\mathbb{R}^{3}$ (22)
@@
PRoPE effectively encodes camera geometry. On both RealEstate10K and Objaverse, we observe that PRoPE is comparable with or outperforms GTA+CamRay, despite containing the same information. PRoPE is also simpler: it is only applied at the attention-level, while GTA+CamRay includes both attention-level and token-level terms. Adding CamRay to PRoPE results in additional improvements to view synthesis quality.
Token-level and attention-level conditioning techniques are complementary. GTA and PRoPE both benefit from the extra CamRay input. GTA benefits significantly more: this can be explained by the fact that the standard SE(3)-based GTA formulation doesn’t encode intrinsics.
4.5 Out-of-distribution Robustness
@@
(a) Testing with {2, 4, 8, 16} input views.
[…]
Figure 4: Evaluation on RealEstate10K [12]. Relative encodings methods demonstrate superior robustness when handling (a) varying numbers of input views and (b) different focal lengths at test time.
@@
One hypothesis for why relative camera encodings outperform absolute ones is improved generalization characteristics; this is similar to why RoPE [28] can improve performance for language modeling. To test this, we benchmark conditioning methods using test-time settings that introduce distribution shifts in sequence length and intrinsics (Figure 3). This lets us evaluate the impact of various camera conditioning techniques on robustness to unseen camera configurations.
Setting 1: Longer Input Sequences at Test Time. Inspired by test-time context length extrapolation [59, 60], our first setting deploys NVS models trained with a fixed number of input views (2 in our experiments) to significantly more views (up to 16) at test time. This is particularly important in real-world scenarios where the number of observations can vary substantially across different applications, and sometimes dynamically increase.
Setting 2: Out-of-distribution Intrinsics at Test Time. Our second setting evaluates a model’s ability to handle varying focal lengths at test time. This is crucial as focal length can vary significantly across different cameras and zoom levels, and it is impractical to train a separate model for every possible focal length. We test our models with focal lengths ranging from 1$\times$ to 5$\times$ the training focal length, simulating scenarios where more zoomed-in images are seen during deployment.
Relative encodings improve generalization; PRoPE outperforms alternatives.
Evaluated results on the RealEstate10K [12] dataset are summarized in Figure 4 with visuals provided in Figure 5 and 6. We make three main observations. First, while Plücker Raymap encodes more complete camera information than CAPE and GTA, it consistently underperforms across all settings—even when intrinsics information is critical. Second, PRoPE demonstrates improved performance and robustness compared to all other methods in both out-of-distribution settings, particularly when handling out-of-distribution focal lengths. This indicates that explicitly modeling the relative projective relationship between cameras is more effective than modeling only the relative SE(3) relationship, as done in GTA and CAPE. Finally, we found adding CamRay to PRoPE actually hurts performance for intrinsics extrapolation; this suggests that PRoPE is uniquely useful for intrinsics extrapolation.
@@ 4.6 Task Generalization
Like our results so far, prior studies on relative pose encodings [3, 4] for multiview transformers have focused experiments on novel view synthesis. To better understand how these conclusions generalize, we evaluate PRoPE in two new tasks: stereo depth estimation using UniMatch [14] and a spatial cognition task designed around DL3DV [15].
@@
RGBD [61] UniMatch 0.123 0.175† 0.678 0.203
@@
Scenes11 [63] UniMatch 0.065 0.085 0.575 0.126
@@
Table 4: Performance Improvement on Stereo Depth Estimation Task with UniMatch [14]. †The “Sq Rel” metric is less reliable on the RGBD dataset because the depth and camera pose are imperfect [63].
Figure 7: Qualitative Results on Stereo Depth Estimation Task. Attention-level camera conditioning in UniMatch [14] leads to significant estimation improvements.
@@
Table 5: Spatial cognition results. We report the accuracy of detecting inconsistent image-camera pairs on the DL3DV [15] dataset under varying numbers of input views. Both CamRay and PRoPE significantly help with performance, without introducing additional model parameters. An illustration of this task can be found in Figure A.3.
Stereo Depth Estimation. We set up this task using UniMatch [14], a pretrained multi-view transformer that has seen wide adoption in downstream applications [64, 65, 66, 67]. UniMatch was originally trained on three different tasks; we focus on the stereo depth estimation task, which assumes known relative camera poses between input views. We incorporate camera information into UniMatch’s cross-view attention mechanism using PRoPE, modifying only $\sim$50 lines of the official code. All models follow the exact same training protocols as described in the original paper.
Spatial Cognition. Next, we design a spatial cognition task inspired by [68]. In contrast to view synthesis and depth estimation where performance hinges on precise geometric understanding for continuous outputs, our spatial cognition task provides a discriminative output for assessing multiview understanding. In this task, the system is given multiple images of the same scene, each paired with camera information. The problem is designed so that it cannot be solved by analyzing the camera information alone, the images alone, or without reasoning about the multiview relationships among all inputs. One of the image-camera pairs is intentionally corrupted by assigning it an incorrect camera pose sampled from other frames. The system is then required to identify the incorrect image-camera pair based on geometric consistency. See Appendix A.1.3 for implementation details, and Figure A.3 for examples of inputs and desired outputs.
@@ 4.7 Scaling PRoPE
Table 6: Scaling LVSM with increased compute. We compare LVSM [8] models trained with Plücker raymaps versus PRoPE on RealEstate10K [12] at two compute scales.
3-view 6-view 9-view
Zip-NeRF [69] 12.77 0.271 0.705 13.61 0.284 0.663 14.30 0.312 0.633
ZeroNeRF [70] 14.44 0.316 0.680 15.51 0.337 0.663 15.99 0.350 0.655
ReconFusion [71] 15.50 0.358 0.585 16.93 0.401 0.544 18.19 0.432 0.511
CAT3D [7] 16.62 0.377 0.515 17.72 0.425 0.482 18.67 0.460 0.460
CAT3D [7] + PRoPE 16.93 0.382 0.505 18.01 0.443 0.479 18.98 0.474 0.461
Table 7: Incorporating PRoPE into CAT3D. Adding PRoPE to the CAT3D [7] multiview diffusion model yields improvements over the original model, with zero additional model parameters and negligible computational overhead.
In our final set of experiments, we evaluate how the advantages of relative camera encoding extends to larger models with more computational resources. We conducted two experiments for this, which are discussed below.
PRoPE’s benefits persist when scaling LVSM. Finally, we scale our LVSM training pipeline with approximately 100$\times$ more computational resources (details listed in Appendix A.1.1). We trained two variants of this larger LVSM model: one that follows the original LVSM paper [8] and uses Plücker rays and one that incorporates PRoPE. Results are reported in Table 6, where we observe that the relative PRoPE encoding continues to improve model quality—by a smaller but still noticeable margin—on larger model variants trained with more resource.
PRoPE improves CAT3D results. With assistance from the original authors, we add PRoPE to and retrain CAT3D [7], a large multiview diffusion model conditioned on näive raymaps. We report metrics from this model in Table 7. PRoPE produces consistent improvements across metrics, while introducing zero additional model parameters and negligible computational overhead.

### v2

@@ 3.4 Projective Position Encoding (PRoPE)
To implement PRoPE, we define a new set of $\mathbf{D}^{\texttt{PRoPE}}_{t}\in\mathbb{R}^{d\times d}$ matrices and use GTA-style attention (Equation 14) to inject them into transformer blocks. We design these matrices to (1) encode frustum relationships between cameras—this uses the projective relationship in Equation 16—and (2) encode relative patch positions within cameras—this follows GTA [miyato2023gta] and uses RoPE terms. These goals are achieved with complementary submatrices, each with shape $\frac{d}{2}\times\frac{d}{2}$:
@@
In these definitions, $\text{RoPE}_{d/4}(\cdot)$ constructs $\frac{d}{4}\times\frac{d}{4}$ rotary embeddings [su2024roformer] for $(x_{t},y_{t})$, which are the patch coordinates for token $t$.
3.5 Properties of PRoPE
@@
Reduction to RoPE [heo2024rotary]. Equation 20 evaluates to identity for patches from the same image. For these token pairs, $\mathbf{D}_{t}^{\texttt{PRoPE}}$ contains only the RoPE terms used by single-image vision transformers.
@@ 4.1 Experiment Setup
We include metrics for several camera conditioning techniques—Naive and Plücker raymap encodings, CAPE [kong2024eschernet], GTA [miyato2023gta], and PRoPE. In our experiments, GTA refers to the SE(3)+SO(2) variation studied by [miyato2023gta], where SO(2) refers to RoPE on patch positions. As discussed in Section 3.4, PRoPE adopts the self-attention mechanism and RoPE combination proposed by GTA. The primary difference between PRoPE and GTA is therefore the use of relative projective relationships instead of relative SE(3) relationships between cameras.
Our core experiments evaluate camera conditioning techniques using feedforward novel view synthesis (NVS). NVS is an ideal benchmarking task because it requires fine-grained geometric reasoning: models are trained to render scenes from target viewpoints, given only calibrated reference images and target camera parameters. We do this by reimplementing and training variants of LVSM [jin2024lvsm], a state-of-the-art novel view synthesis method that originally encoded camera geometry with Plücker raymaps. We train and evaluate separately on the RealEstate10K [zhou2018stereo] and Objaverse [deitke2023objaverse] datasets. Scenes in RealEstate10K are captured with constant intrinsics, but cameras vary between scenes. Objaverse renders use the same camera intrinsics across the entire dataset.
We take several steps to ensure fairness in evaluations. Models are trained in the same codebase, with matched hyperparameters and training steps. All main experiments use identical input, output, and overall model sizes ($\sim$25M parameters); we also validate larger models in Section 4.7. More details are provided in Appendix A.1.1.
@@ 4.2 Relative vs Absolute Positional Encodings
Method RealEstate10K [zhou2018stereo] Objaverse [deitke2023objaverse]
@@
Naive Raymap 20.54 0.210 0.623 21.59 0.153 0.856
CAPE [kong2024eschernet] 21.11 0.234 0.656 19.68 0.220 0.827
GTA [miyato2023gta] 22.51 0.164 0.707 23.70 0.104 0.879
@@
Table 1: Novel view synthesis comparison, with constant intrinsics in each scene. We compare different camera conditioning approaches applied to the LVSM [jin2024lvsm] framework.
Novel view synthesis results are presented in Table 1 and discussed below.
Relative encodings outperform absolute ones. Consistent with prior work [kong2024eschernet, miyato2023gta], we observe that relative encodings for camera geometry consistently outperform absolute ones. CAPE, GTA, and PRoPE all yield improvements over widely used raymap encodings, with PRoPE (followed closely by GTA) producing the best results.
Projective positional encoding improves view synthesis quality. PRoPE consistently outperforms other encoding methods across metrics on RealEstate10K [zhou2018stereo], despite the dataset’s limited intrinsics variation. This confirms that capturing more complete camera information (both intrinsic and extrinsic) in our relative encoding is beneficial. We also find no loss of performance when the train and test images all have constant intrinsics: GTA [miyato2023gta] and PRoPE produce identical metrics for the Objaverse dataset, which verifies that PRoPE reduces to GTA when camera intrinsics are unimportant.
Method RealEstate10K [zhou2018stereo] Objaverse [deitke2023objaverse]
@@
Naive Raymap 20.56 0.301 0.629 21.00 0.191 0.845
CAPE [kong2024eschernet] 15.94 0.497 0.699 16.78 0.322 0.760
GTA [miyato2023gta] 15.77 0.512 0.641 18.00 0.257 0.775
@@
Table 2: Novel view synthesis, with varying intrinsics in each scene. We compare the LVSM [jin2024lvsm] model trained with different camera conditioning strategies, on intrinsics-augmented dataset variants.
@@ 4.3 Attention-Level Intrinsics Conditioning
Real-world data often involves different cameras and focal lengths—consider multiview rigs on autonomous cars or zoom lenses on point-and-shoot cameras. To understand how effective PRoPE is at encoding intrinsics information, we evaluate each conditioning method on intrinsics-augmented versions of the RealEstate10K [zhou2018stereo] and Objaverse [deitke2023objaverse] datasets. We augment RealEstate10K by applying a zoom factor sampled uniformly in $[1,3]$ to each image. For Objaverse, we switch from constant field of views to uniformly sampled ones between 35 and 50 degrees. In contrast to Section 4.2, this means that cameras within scenes can vary in both extrinsics and intrinsics. We present quantitative results in Table 2 and qualitative results in Figure A.1.
PRoPE enables intrinsic-aware multiview understanding. We observe that PRoPE outperforms all alternative camera conditioning techniques for both the RealEstate10K and Objaverse datasets. Existing attention-based methods, which let networks condition on relative pose, perform poorly without knowledge of intrinsics. While token-level raymaps carry sufficient camera information, they perform worse overall than PRoPE’s relative conditioning formulation.
@@ 4.4 Hybrid Encoding Strategies
Figure 2: Hybrid camera encoding. Left: token-level conditioning using only raymaps, which capture both intrinsics and extrinsics. Right: hybrid encoding where attention-level PRoPE captures relationships between camera frustums, and token-level CamRay encodes local ray directions.
Method RealEstate10K [zhou2018stereo] Objaverse [deitke2023objaverse]
@@
Table 3: Novel view synthesis with hybrid camera encodings, with varying intrinsics in each scene. Intrinsics can be conditioned by concatenating local frame camera rays to the network input.
Token- and attention-level camera encodings require modifications to different parts of the transformer architecture. They are therefore compatible with each other: both conditioning styles can be used simultaneously. To compare PRoPE against a stronger baseline while simultaneously exploring these “hybrid” conditioning strategies (Figure 2), we train LVSM variations that couple relative encodings with local, camera-frame raymaps:
$\mathbf{M}_{i,\texttt{CamRay{}}}^{u,v}$ $=\mathbf{R}^{\text{cw}}_{i}\mathbf{d}_{i}^{u,v}\propto\boldsymbol{K}_{i}^{-1}\begin{bmatrix}u&v&1\end{bmatrix}^{\top}\in\mathbb{R}^{3}$ (21)
@@
PRoPE effectively encodes camera geometry. On both RealEstate10K and Objaverse, we observe that PRoPE is comparable with or outperforms GTA+CamRay. This is true even though PRoPE is simpler: it is only applied attention-level, while GTA+CamRay includes both attention-level and token-level terms.
Token-level and attention-level conditioning techniques are complementary. GTA and PRoPE both benefit from the extra CamRay input. GTA benefits significantly more; this can be explained by the absence of intrinsics information in the standard SE(3)-based GTA formulation.
@@
((a))
((b))
Figure 4: Evaluation on RealEstate10K [zhou2018stereo]. Relative encoding methods demonstrate superior robustness on handling (a) varying numbers of input views and (b) different focal lengths at test time.
@@
4.5 Out-of-distribution Robustness
One hypothesis for why relative camera encodings outperform absolute ones is improved generalization characteristics; this is similar to why RoPE [heo2024rotary] can improve performance for language modeling. To test this, we benchmark conditioning methods using test-time settings that introduce distribution shifts in sequence length and intrinsics (Figure 3).
Setting 1: Longer Input Sequences at Test Time. Inspired by test-time context length extrapolation [pope2023efficiently, wang2024beyond], our first setting deploys NVS models trained with a fixed number of input views (2 in our experiments) to significantly more views (up to 16) at test time. This is particularly important in real-world scenarios where the number of observations can vary substantially across different applications, and sometimes dynamically increase.
[…]
Relative encodings improve generalization; PRoPE outperforms alternatives. Evaluated results on the RealEstate10K [zhou2018stereo] dataset are summarized in Figure 4, with visuals provided in Figure 5 and 6. We make three main observations. First, while Plücker Raymap encodes more complete camera information than CAPE and GTA, it consistently underperforms across all settings—even when intrinsics information is critical. Second, PRoPE improves performance and robustness in both out-of-distribution settings, particularly when handling out-of-distribution focal lengths. This indicates that explicitly modeling the relative projective relationship between cameras is more effective than modeling only the relative SE(3) relationship, as done in GTA and CAPE. Finally, we found adding CamRay to PRoPE actually hurts performance for intrinsics extrapolation; this suggests that PRoPE is uniquely useful for intrinsics generalization.
@@ 4.6 Task Generalization
Like our results so far, prior studies on relative pose encodings [miyato2023gta, kong2024eschernet] for multiview transformers have focused experiments on novel view synthesis. To better understand how these conclusions generalize, we evaluate PRoPE in two new tasks: stereo depth estimation using UniMatch [xu2023unifying] and a spatial cognition task designed around DL3DV [ling2024dl3dv].
@@
RGBD [sturm2012benchmark] UniMatch 0.123 0.175† 0.678 0.203
@@
Scenes11 [ummenhofer2017demon] UniMatch 0.065 0.085 0.575 0.126
@@
Table 4: Performance Improvement on Stereo Depth Estimation Task with UniMatch [xu2023unifying]. †The “Sq Rel” metric is less reliable on the RGBD dataset due to the imperfect depth and camera pose [ummenhofer2017demon].
Figure 7: Qualitative Results on Stereo Depth Estimation Task. Attention-level camera conditioning in UniMatch [xu2023unifying] leads to significant estimation improvements.
@@
Table 5: Spatial cognition results. We report the accuracy of detecting inconsistent image-camera pairs on the DL3DV [ling2024dl3dv] dataset under varying numbers of input views. Both CamRay and PRoPE significantly help with performance, without introducing additional model parameters. An illustration of this task can be found in Figure A.3.
Stereo Depth Estimation. We set up this task using UniMatch [xu2023unifying], a pretrained multiview transformer that has seen wide adoption in downstream applications [chen2024mvsplat, chen2025mvsplat360, smith2024flowmap, niu2024mofa]. UniMatch was originally trained on three different tasks; we focus on the stereo depth estimation task, which assumes known relative camera poses between input views. We incorporate camera information into UniMatch’s cross-view attention mechanism using PRoPE, modifying only $\sim$50 lines of the official code. All models follow the exact same training protocols as described in the original paper.
Spatial Cognition. Next, we design a spatial cognition task inspired by [bonnen2025evaluating]. In this task, a network is given multiple images of the same scene, each paired with camera information. The problem is designed so that it cannot be solved by analyzing the camera information alone, the images alone, or without reasoning about the multiview relationships among all inputs. One of the image-camera pairs is intentionally corrupted by assigning it an incorrect camera pose sampled from other frames. The network is then required to identify the incorrect image-camera pair based on geometric consistency. See Appendix A.1.3 for implementation details, and Figure A.3 for input-output examples.
@@ 4.7 Scaling PRoPE
In our final set of experiments, we evaluate how the advantages of relative camera encoding extend to larger models with more computational resources. We conducted two experiments for this, which are discussed below.
[…]
PRoPE improves CAT3D results. With assistance from the original authors, we add PRoPE to and retrain CAT3D [gao2024cat3d], a large multiview diffusion model conditioned on naive raymaps. We report metrics from this model in Table A.2. PRoPE produces consistent improvements across metrics, while introducing zero additional model parameters and negligible computational overhead.
@@
Table 6: Scaling LVSM with increased compute. We compare LVSM [jin2024lvsm] models trained with Plücker raymaps versus PRoPE on RealEstate10K [zhou2018stereo] at two compute scales.